OpenStudy (anonymous):

LIL BROS MATHS: I NEED HELP. Use algebra to determine the dimensions of a spare based popper box that will hold 250ml, yet has the smallest possible surface area. Use a graphics calculator and show all working out........anyone know how to do this properly? There are lots of answers.....and no google help..

OpenStudy (anonymous):

I am going to try to solve this by hand.

OpenStudy (anonymous):

haha ok....i need epic help. :-(

OpenStudy (anonymous):

square box of vol 250 ml minimize: smallest possible surface area requires light algebra will attempt in morning if no one gets back to you

OpenStudy (anonymous):

If the box is square shaped, all its sides are the same length. So length=width=height volume=length x width x height. 250=x^3 x=6.30=length=width=height.

OpenStudy (anonymous):

unless i interpreted this question incorrectly and "spare based popper box" does not refer to a cube.

OpenStudy (anonymous):

the box isn't square shaped,it's a rectangular prism (base is square+top,but 4 faces are rectangular)

OpenStudy (anonymous):

OK \[V =250=x ^{2}y\]\[y =250/x ^{2}\]\[SA =2xy+2xy+2x ^{2}=4xy+2x ^{2}\]Find derivative of SA. Minimise